OpenStudy (anonymous):

What does it mean for a number to be the "solution" to an equation? Give an example of a variable equation and its solution. Explain in complete sentences how you know that number is the solution.

Parth (parthkohli):

If a given quantity is a solution to an equation, it means that the LHS and RHS will be equal when the quantity is plugged in for the variable.

OpenStudy (anonymous):

I am in the beginning of algebra one 9th grade i didnt understand what they are asking for

OpenStudy (anonymous):

Hey All !

OpenStudy (anonymous):

So , Mr. @Coyzita what is NOT an equation ( first let us understanding on that) :

Parth (parthkohli):

For finding the solution, you have to solve for the variable.

OpenStudy (anonymous):

3= 16-13 is NOT an equation because it is a mathematical relation that is just true. ALWAYS TRUE

OpenStudy (anonymous):

3 = 16 -12 is Also NOT an equat. because it is ALWAYS FALSE

OpenStudy (anonymous):

Now, we can UNITE BOTH of these situation in a mathematical relation 3 = 16 - X This time, THIS RELATION is SOMETIMES TRUE, and SOMETIMES FALSE

OpenStudy (anonymous):

"What do I mean sometimes ?" :

OpenStudy (anonymous):

One can PUT A NUMBER VALUE INSTEAD OF THE LETTER X. We cal it "substitute" or less formally "plug in"

OpenStudy (anonymous):

There will be some values , which after "plugging in" will give REALLY TRUE EQUALITY. Such numbers are called SOLUTIONS TO THE EQUATION

OpenStudy (anonymous):

Repeat: 3 = 16 - 12 is NOT an EQUATION. It is TRUE ALWAYS. 3 = 16 - X is an EQUATION. The solution is a number that after being SUBSTITUTED instead of the letter X will give TRUE EQUALITY

OpenStudy (anonymous):

So - an expression containing Variable letters (Variables ) and two sides with required equality sign between them is an equation. Solution is the value that one can plug instead of the letter and get TRUE expression

OpenStudy (anonymous):

Mister @Coyzita ?

OpenStudy (anonymous):

gone he is, isn't he , hmmm .....